OpenStudy (anikate):

Write a rule for the nth term of the geometric sequence. 5, 10, 20, 40, ...

OpenStudy (anikate):

@TheViper

OpenStudy (anonymous):

n=2(n-1) i want to say thats it

OpenStudy (anikate):

then again how will that help?

OpenStudy (anikate):

@NickR what do i substitute

OpenStudy (anonymous):

\[a_{n}=5*2^{n-1}\]

OpenStudy (anikate):

what's that?

OpenStudy (anikate):

what's the actual formula?

OpenStudy (anikate):

@myko

OpenStudy (anonymous):

ah forgot the 5

OpenStudy (anonymous):

thats your answer

OpenStudy (anikate):

can someone plz just tell me the formula and what to substitute

OpenStudy (anonymous):

myko gave you the answer

OpenStudy (anikate):

is that the answer or formula?

OpenStudy (anonymous):

that is your rule for the function

OpenStudy (anikate):

how did he find it?

OpenStudy (anikate):

what's the formula?

OpenStudy (anonymous):

well you can see that its exponential, and is 2* the number before it, and starts at 5 so...

OpenStudy (anonymous):

there really isnt a general formula

OpenStudy (anikate):

plz just explain to me again how he got the answer ? also what's the answer?

OpenStudy (anikate):

@NickR

OpenStudy (anonymous):

you can see that the sequence is exponential right

OpenStudy (anikate):

exponential? what's that?

OpenStudy (anonymous):

means that it doesnt increase at a constant rate

OpenStudy (anikate):

alright. in this sequence the number is multiplied by 2to get the next number.

OpenStudy (anonymous):

correct

OpenStudy (anikate):

right?

OpenStudy (anonymous):

you got it because i need to go

OpenStudy (anikate):

what do i have to do next?

OpenStudy (anonymous):

and the initial number is 5 so its 5*2(n-1) the general equation for this equation is \[a _{n}=a _{1}*d(n-1)\]

OpenStudy (anonymous):

where a1 is your first term 5 and d is how many times its multiplied 2

OpenStudy (anikate):

so its 10(n-1) as the answer

OpenStudy (anikate):

@NickR

OpenStudy (anikate):

@satellite73

OpenStudy (anonymous):

\[a_n=5(2)^{n-1}\]

OpenStudy (anonymous):

\[a_n=a(r)^{n-1}\] a=first term r=ratio